OpenStudy (kathydimieri12):

 Question 3 (Multiple Choice Worth 5 points) (05.01)A window is 4 feet wide and 8 feet tall. A rectangle is shown. The length of the rectangle is labeled 4 feet. The width of the rectangle is labeled 8 feet. A manufacturer wants to use a scale factor of 1.5 to enlarge a window. What will the area of the window be after it is enlarged? 32 ft2 48 ft2 52.25 ft2 72 ft2

OpenStudy (er.mohd.amir):

new length=1.5*4=6 new breath=1.5*8 =12 area=new length*new breath=6*12=72 feet

OpenStudy (kathydimieri12):

thank you